### v1

# Quantum Critical Collapse Abhors a Naked Singularity

Abstract
Classical critical collapse yields naked singularities from smooth initial data, challenging cosmic censorship and shaping the spectrum of primordial black holes. We show that one-loop vacuum polarization near the threshold alters this outcome. In analytically tractable Einstein-scalar critical spacetimes, regularity uniquely selects a Boulware-like state whose stress tensor supplies a universal quantum growing mode. Its backreaction competes with the classical unstable mode, producing a shift of the critical point and a finite mass gap at the new threshold, thereby enforcing horizon formation even under arbitrary fine-tuning. In primordial collapse, the threshold shift enters exponentially into the formation fraction, while the gap truncates the low-mass tail—effects that may reshape the predicted mass spectrum. These results provide the first consistent quantum treatment of critical collapse, offering definitive predictions for several long-standing problems.
The fate of spacetime singularities is of central importance in gravity: every classical black hole culminates in one, and every expanding universe originates from one. If such singularities are not hidden behind horizons, classical determinism fails, compelling us to confront quantum gravity. At the heart of this conundrum lies the cosmic censorship conjecture [1]; as Hawking remarked, “Nature abhors a naked singularity” [2].
However, explicit counterexamples involving naked singularities are known. The most infamous one is the endpoint of black hole evaporation, where Hawking radiation shrinks the horizon and may ultimately reveal a singularity [3, 4]. Another notable example is the Gregory-Laflamme instability [5], in which higher-dimensional black strings undergo classical fragmentation, possibly producing naked pinch-off singularities.
Yet the most physically realistic setting is classical critical gravitational collapse [6, 7], wherein finely tuned smooth initial data drive the formation of naked singularities. This process can arise in the early universe and has been identified as a universal feature in scenarios involving primordial black hole (PBH) formation [8]. In this Letter, we show that quantum effects fundamentally overturn the classical picture: one-loop vacuum polarization generates a universal growing mode that enforces horizon formation and prevents naked singularities, even under arbitrarily fine-tuned initial data. This provides, for the first time, a consistent semiclassical description of critical collapse.
Classical critical collapse.
Critical collapse was first discovered by Choptuik [6] in the spherically symmetric collapse of a massless scalar field—the prototypical model initiated by Christodoulou, referred to as the Einstein-scalar system [9, 10, 11, 12, 13]. It probes the threshold of black hole formation in the space of initial data. By varying any one-parameter family $p$ of initial configurations between dispersion and black hole formation, a critical value $p^{\ast}$ emerges. For marginally supercritical data $p>p^{\ast}$, the black hole mass follows a universal scaling law:
$M_{\text{BH}}\propto(p-p^{\ast})^{\gamma},$ (1)
where the exponent $\gamma=0.37$ remarkably depends only on the type of collapsing matter. As $p\to p^{\ast}$, a zero-mass black hole forms, yielding a naked singularity. By analogy with statistical mechanics, $M_{\text{BH}}$ functions as an order parameter, distinguishing Type I and Type II collapses by the presence or absence of a mass gap. The Einstein-scalar system exhibits robust, universal Type II critical behavior.
In the high-curvature region preceding black hole formation, the spacetime evolves toward a self-similar critical solution, independent of the initial data family [14, 15, 16, 17, 18]. Self-similarity refers to the fact that the solution repeats its structure at progressively smaller scales, either continuously or discretely. Choptuik’s critical solution exhibits discrete self-similarity (DSS), where the solution recurs only after a fixed logarithmic rescaling known as the echoing period.
On the other hand, continuously self-similar (CSS) critical solutions are analytically tractable and admit a homothetic Killing vector
$\mathcal{L}_{\xi}g_{\mu\nu}=2g_{\mu\nu},\quad\xi=-\frac{\partial}{\partial T},$ (2)
allowing the metric to be expressed as
$g_{\mu\nu}(T,x^{i})=\ell^{2}e^{-2T}\tilde{g}_{\mu\nu}(x^{i}),$ (3)
where the length scale $\ell$ is arbitrary. In the critical spacetime, curvature invariants diverge as $T\to\infty$, corresponding to the naked singularity (see Fig. 1). The solution acts as a codimension-one attractor in phase space, characterized by exactly one unstable spherically symmetric mode [14, 15]. All other perturbations decay [19], making the solution unstable in the mildest possible way. Physically, this single growing mode determines whether the collapse evolves toward black hole formation or dispersion.
Figure 1: Global structure of a CSS critical spacetime, in which the past light cone of the naked singularity naturally separates the spacetime into interior and exterior regions [16, 17, 20]. Our focus is on the interior region, where the singularity dynamically emerges from smooth initial data.
Quantum critical collapse.
The classical picture is clearly incomplete. In the self-similar phase of Type II collapse, curvature grows without bound on ever-diminishing scales, precisely where quantum effects should start to play a role. This regime offers an ideal laboratory for incorporating quantum corrections from collapsing matter, long before a full quantum gravity treatment becomes unavoidable. The question we address is whether quantum effects step in, protecting the singularities from being naked in the end.
Although the issue is fundamental and the expectation fairly intuitive, semiclassical analyses of critical spacetimes have received surprisingly little attention over the past thirty years [21, 22, 23, 24, 25, 26, 27, 28], and a consistent, definitive treatment remains elusive. Most approaches rely on conformal matter or special quantum states, assumptions that miss essential aspects of the critical dynamics or introduce spurious Hawking flux even before black hole formation. More seriously, these methods have produced widely divergent conclusions regarding the role of quantum effects in critical collapse [29].
The root of these difficulties is simple: quantum corrections are notoriously challenging to handle in genuinely dynamical curved spacetimes, especially without conformal matter. We aim to overcome these obstacles, establishing a rigorous, first-principles semiclassical framework for critical collapse in the Einstein-scalar system.
Since all non-spherical modes decay except a single spherical growing mode, it is natural to focus exclusively on the $s$-wave sector, which captures the dominant quantum corrections. For a minimally coupled, massless scalar field $f$ in $D=d+1$ dimensions, one isolates the $s$-wave via spherical reduction to two dimensions with a dilaton $\phi$ [30]:
$ds^{2}_{(D)}=g_{ab}dx^{a}dx^{b}+e^{\frac{-4\phi}{D-2}}d\Omega^{2}_{D-2},$ (4)
under which the matter Lagrangian acquires a universal dilaton coupling
$\mathcal{L}_{\text{matter}}\propto-\sqrt{-g}e^{-2\phi}(\nabla f)^{2}.$ (5)
Although the reduced matter sector is no longer a free conformal theory, it retains two-dimensional Weyl invariance. This property facilitates a well-defined computation of the one-loop exact trace anomaly via the heat kernel method in path integral quantization, shown to be independent of both the regularization scheme and the choice of quantum state. The trace of the renormalized stress tensor is given by [31, 32, 33, 34, 35, 36, 37]
$\langle T^{a}{}_{a}\rangle=\frac{\hbar}{24\pi}(R-6(\nabla\phi)^{2}+6\Box\phi).$ (6)
Here $R$ reproduces the conformal anomaly of a free scalar, while the dilaton terms encode the effects of spherical reduction, ensuring that the essential quantum imprint of the higher-dimensional theory is captured. Crucially, the universal dilaton coupling makes the anomaly structure extremely general across different matter models.
The one-loop effective action contains nonlocal anomaly-induced pieces whose variation yields Eq. (6):
$\Gamma\sim-\frac{\hbar}{96\pi}\int(R\Box^{-1}R-12(\nabla\phi)^{2}\Box^{-1}R+12\phi R).$ (7)
Additional Weyl-invariant terms may also appear. While they do not affect the trace, they typically lack closed-form expressions and can be sensitive to the chosen quantum state [38, 39, 40]. This has been a known source of confusion in black hole evaporation [41, 42, 43, 44, 45, 37, 46], which was clarified in [47]. Fortunately, the regular state appropriate to critical collapse is insensitive to these subtleties [29].
To render the one-loop action local, we introduce auxiliary fields $\chi_{1}$ and $\chi_{2}$ satisfying
$\Box\chi_{1}=\lambda_{1}R+\lambda_{2}\left(\nabla\phi\right)^{2},\quad\Box\chi_{2}=-\mu_{1}R-\mu_{2}\left(\nabla\phi\right)^{2},$ (8)
with coefficients $\{\lambda_{i},\mu_{i}\}$ determined by matching to Eq. (7). The homogeneous solutions for $\chi_{1,2}$ encode the boundary conditions, thereby specifying the quantum state. Varying the one-loop action with respect to the metric determines $\langle T_{ab}\rangle$ [29].
Finally, compatibility with the conservation law lifts this result back to the full $D$-dimensional stress tensor via the $s$-wave relations [31]
$\langle T^{(D)}_{ab}\rangle\propto\frac{\langle T^{(2)}_{ab}\rangle}{e^{-2\phi}},\quad\langle T^{(D)}_{\theta\theta}\rangle\propto e^{\frac{2\phi(D-4)}{D-2}}\frac{1}{\sqrt{-g}}\frac{\delta\Gamma}{\delta\phi}.$ (9)
The anomaly-based computation of one-loop backreaction is robust, insensitive to ambiguities that generically arise in curved spacetimes, and consistent with Wald’s axiomatic framework [48, 49, 50]. To demonstrate the power and generality of our formalism, we focus on two exact CSS critical solutions available in closed form for the Einstein-scalar system—the Garfinkle spacetime in $D=2+1$ [51] and the Roberts spacetime in $D=3+1$ [52, 53, 54]. As demonstrated in various studies [55, 56, 57, 58, 59, 60], these solutions do not meet the strict criteria of critical spacetimes but closely mirror numerically determined critical behavior and capture their salient features. The conclusions we draw are not confined to these particular solutions, but are generic to self-similar critical spacetimes.
Garfinkle spacetime in 2+1 dimensions.
This family of critical spacetimes is labeled by a positive integer $n$, with the metric:
$ds^{2}=e^{-2T}\bigg{[}e^{2\rho_{0}}\bigg{(}dx-\frac{x}{2n}dT\bigg{)}dT+r_{0}^{2}d\theta^{2}\bigg{]},$ (10)
where
$e^{2\rho_{0}}=2n\bigg{(}\frac{1+x^{n}}{2}\bigg{)}^{4(1-\frac{1}{2n})},\quad r_{0}=\frac{1-x^{2n}}{2},$ (11)
with $T\in(-\infty,\infty)$ and $x\in[0,1]$. The scalar field supporting this geometry is (with units $c=1$, $8\pi G_{N}=1$)
$f(T,x)=\sqrt{\frac{2n-1}{2n}}\bigg{[}T-2\ln{\bigg{(}\frac{1+x^{n}}{2}\bigg{)}}\bigg{]}.$ (12)
A curvature singularity appears as $T\to\infty$, with the global structure resembling the interior region of Fig. 1. In particular, the $n=4$ case shows remarkable agreement with numerical simulations [56].
The analytic structure of this geometry permits closed-form expressions for $\langle T^{(3)}_{\mu\nu}\rangle$ of any integer $n$. Surprisingly, regularity within the self-similar domain $x\in[0,1]$ fixes the homogeneous parts of the auxiliary fields and uniquely selects a Boulware‐like vacuum state. The state is manifestly asymptotically Minkowskian and features a time-independent two-dimensional $\langle T_{ab}\rangle$. Only genuine vacuum polarization contributes, with no spurious radiation at infinity. One finds the particularly simple form:
$\langle T^{(3)}_{\mu\nu}\rangle=e^{T}F_{\mu\nu}(x,n),$ (13)
where $F_{\mu\nu}(x,n)$ is real-analytic in $x$, available in [29]. The overall $e^{T}$ growth has a clear physical origin: it reflects the self‐similar rescaling of the areal radius, carried by the dilaton $\phi$ through the $s$-wave reduction in Eq. (9). The one-loop quantum corrections introduce a genuine growing mode and are thus indispensable.
A full treatment must also include the classical growing mode $(p-p^{\ast})e^{\omega_{c}T}$, arising from deviations off the critical point, where the exponent $\omega_{c}$ for each $n$ has been worked out in [55, 56]. This allows us to study the competition between quantum vacuum polarization and classical supercritical evolution in shaping the near-critical collapse.
(a) $n=4$ Garfinkle spacetime.
(b) Roberts spacetime.
Figure 2: (a) For the most physically relevant $n=4$ Garfinkle solution, we plot the logarithm of $M_{\text{EMOTS}}$ against the ratio defined in Eq. (17). As $\mathcal{R}$ decreases, effectively moving toward $p\to p^{\ast}_{q}$, the mass function monotonically approaches a constant value, indicating a Type I behavior. (b) For the Roberts spacetime, note the Hawking mass in four dimensions is defined as $M\equiv\frac{\bar{r}}{2}[1-(\nabla\bar{r})^{2}]$. As $\mathcal{R}$ decreases, quantum corrections cause an abrupt, nonmonotonic change in the behavior of $M_{\text{EMOTS}}$.
The backreacted geometry then acquires quasi-CSS perturbations of the form
$e^{2\rho(T,x)}$ $=$ $e^{2\rho_{0}}+(p-p^{\ast})F_{c}(x)e^{\omega_{c}T}+\hbar F_{q}(x)e^{\omega_{q}T},\quad$ (14)
$r(T,x)$ $=$ $r_{0}+(p-p^{\ast})r_{c}(x)e^{\omega_{c}T}+\hbar r_{q}(x)e^{\omega_{q}T},$ (15)
where $\omega_{q}=1$. For each $n$, the functions $F_{i}(x)$ and $r_{i}(x)$ admit closed, real-analytic forms [29].
To probe horizon formation, we compute the quasi‐local Hawking mass [61, 62, 63, 64]
$M(T,x)\equiv\bar{r}^{2}-(\nabla\bar{r})^{2},\quad\bar{r}=e^{-T}r(T,x).$ (16)
Under spherical symmetry, apparent horizons satisfy $(\nabla\bar{r})^{2}=0$, while $(\nabla\bar{r})^{2}<0$ identifies trapped regions. The nonlinearity of $(\nabla\bar{r})^{2}$ allows perturbations to reach $O(1)$ while remaining within linear perturbation of the background. A careful numerical analysis [29] reveals a surprising feature: quantum corrections lower the critical point from $p^{\ast}$ to a new value $p^{\ast}_{q}<p^{\ast}$. We then quantify the competition between classical and quantum modes by the ratio
$\mathcal{R}\equiv\frac{e^{\omega_{c}T}(p-p^{\ast}_{q})}{e^{\omega_{q}T}\hbar},$ (17)
and evaluate the logarithm of the mass at the earliest marginally outer-trapped surface (EMOTS) where $M_{\text{EMOTS}}=\bar{r}^{2}(T_{\text{EMOTS}},x_{\text{EMOTS}})$. As $p\to p^{\ast}_{q}$, the would‐be Type II collapse acquires a nonzero black hole mass, indicating a phase transition to a quantum-induced Type I behavior with a finite mass gap at the shifted threshold. This is illustrated in Fig. 2a.
Roberts spacetime in 3+1 dimensions.
This analytic CSS spacetime is closely related to the DSS solution first identified by Choptuik [57, 58, 59, 60]. It takes a simple form featuring a null singularity as $T\to\infty$, distinct from the Garfinkle case:
$ds^{2}=2e^{-2T}e^{2x}[(1-e^{-2x})dT^{2}-2dTdx+d\Omega^{2}],$ (18)
with the scalar field $f=\sqrt{2}x$ and $x\in[0,\infty)$. Regularity again fixes a unique vacuum state encoding vacuum polarization, and the resulting stress tensor behaves as [29]
$\langle T^{(4)}_{\mu\nu}\rangle=e^{2T}F_{\mu\nu}(x),$ (19)
with a scaling exponent differing from Eq. (13). A similar backreaction and horizon-tracing analysis reveals a qualitatively similar, but much more exotic transition, as shown in Fig. 2b.
Discussion.
The semiclassical framework based on trace anomaly for critical collapse in the Einstein-scalar system accommodates non-conformal matter and genuinely time-dependent, self-similar geometries without an exact timelike Killing vector. It yields sharp answers to long-standing questions about quantum effects in critical collapse while allowing us to extract general lessons beyond the specific spacetimes studied, which we shall elaborate below.
First, regularity in the critical background unambiguously singles out a Boulware-like vacuum that is asymptotically Minkowskian, with no artificial quantum flux near infinity. Mathematically, such uniqueness is remarkable, since a curved spacetime typically admits infinitely many inequivalent vacua. Physically, however, this is the most reasonable scenario, as it captures the self-energy of collapsing matter before any horizon forms.
Furthermore, the quantum state features a stationary effective two-dimensional stress tensor, yet the physical higher-dimensional ones inherit apparent time dependence through the dilaton-encoded areal radius from Eq. (9). By self-similarity, this yields a universal growing mode $e^{(D-2)T}$, independent of the matter model, for any spherical self-similar critical spacetime.
We therefore regard the existence of a unique Boulware-like state as the a priori physical choice and a generic feature of critical collapse. Our results then imply that vacuum polarization dynamically enforces cosmic censorship: even with arbitrary fine-tuning, would-be naked singularities are shielded by a horizon.
By the same reasoning, we conjecture that all higher-loop corrections must share this universal factor while scaling with additional powers of $\hbar$, for the simple fact that the growing mode is tied to the self-similarity of the background, rather than the loop-counting effect. Schematically,
$\langle T^{(D)}_{\mu\nu}\rangle\propto e^{(D-2)T}(\hbar F_{\mu\nu}(x^{i})+\hbar^{2}\tilde{F}_{\mu\nu}(x^{i})+\cdots).$ (20)
Since higher loops remain parametrically suppressed, this not only strengthens the validity of the linear perturbation analysis but also makes it possible to explore the nonlinear regime nonperturbatively—where both the classical and one-loop quantum modes reach O(1)O(1)—while still remaining within the semiclassical framework.
These conclusions are extracted through analytically tractable CSS critical spacetimes; yet for $D\geq 4$, the true critical spacetimes of the Einstein-scalar system are DSS [7]. Our scaling arguments still apply, where CSS growth $e^{(D-2)T}$ is modulated by a bounded periodic function of the echoing period. This periodic modulation does not enhance the net growth, so the competition between the classical unstable mode and the quantum growing mode, and the threshold shift and mass gap, persist in the DSS case.
Second, the threshold shift and the quantum-induced mass gap indeed mirror classical Type I systems (e.g., massive-scalar and Einstein-Yang-Mills collapse [65, 66]), when a dynamically relevant scale breaks exact self-similarity. The black hole mass near threshold plateaus at a finite universal value, independent of the initial data.
Physically, quantum vacuum polarization furnishes the new scale: $\hbar$ introduces a genuine growing mode that competes with the classical one until nonlinear effects dominate. Because vacuum polarization is generic to quantum matter, the new scale it provides is universally present in all matter systems; accordingly, it is natural to expect that any Type II collapse may generically exhibit Type I behavior once quantum effects are included: intuitively, quantum corrections tend to shield the would-be singularity and facilitate horizon formation, thereby lowering the critical threshold [29].
However, unlike classical Type I, where a metastable soliton (e.g., a boson star or Bartnik-McKinnon solution [67, 68]) with a universal lifetime scaling precedes horizon formation, our framework quantizes matter on a fixed background and treats backreaction linearly. Fully capturing a potential quantum soliton phase will therefore require going beyond the linear regime and implementing dynamical semiclassical simulations—which, given the parametric suppression of higher loops, should be feasible. Encouragingly, recent soliton-like geometries supported by Boulware-like vacua [69, 70, 71, 72, 73, 74, 75, 76, 77, 47, 78, 79] hint that such a nonperturbative phase may indeed exist and may already have been realized. Identifying the requisite metastable phase remains a crucial open question.
Finally, our proposed mechanism may have testable consequences for PBH formation. In the radiation-dominated era modeled as a perfect fluid, the control parameter is the local density contrast at horizon entry $\delta=\frac{\rho-\bar{\rho}}{\bar{\rho}}$, where $\bar{\rho}$ is the mean energy density. There exists a critical threshold $\delta_{c}$ at which PBHs form with a Type II spectrum [8]
$M_{\text{PBH}}=kM_{\text{H}}(\delta-\delta_{c})^{\gamma},$ (21)
where $M_{\text{H}}$ is the Hubble horizon mass. The critical solution for perfect-fluid matter is precisely of CSS type [80, 81, 82], with $\gamma\simeq 0.36$. For a Gaussian density perturbation with rms amplitude $\sigma$, the PBH formation fraction is
$\beta\equiv\frac{\rho_{\text{PBH}}}{\rho_{\text{total}}}\propto\exp{\bigg{(}-\frac{\delta^{2}_{c}}{2\sigma^{2}}\bigg{)}},\quad\sigma\ll\delta_{c}.$ (22)
So even a tiny shift in $\delta_{c}$ translates into exponentially significant changes in abundance. A well-known drawback of PBH scenarios is that observationally compatible models require delicate fine-tuning of $\delta_{c}/\sigma$ [83].
Quantum vacuum polarization—similarly captured for a radiation fluid by the conformal anomaly—adds a second, quantum growing mode that modifies this picture in two model-robust ways: it shifts the threshold by $\Delta\delta$ and produces a universal mass gap $M_{\text{gap}}$. In our scale-free Einstein-scalar system, $M_{\text{gap}}\propto\sqrt{\hbar}$, so one might expect the resulting mass gap to lie near the Planck scale, rendering the semiclassical analysis questionable and seemingly irrelevant for cosmology. Our analysis shows this need not be the case [29]: the proportionality coefficient is sensitive to the magnitudes of the perturbation profiles. In realistic scenarios, these strengths are set, within the regime of semiclassical validity, by a separation-of-scales matching between the near-critical patch and the FRW background [8].
The threshold shift turns out to be epoch-independent [29], and it feeds exponentially into $\beta$ via the relation above, potentially easing fine-tuning. The mass gap converts the classical Type II scaling into a Type I behavior with a universal floor, reshaping the mass spectrum by concentrating PBHs near $M_{\text{gap}}$ rather than at an $O(1)$ fraction of $M_{\text{H}}$ [29, 8, 84, 85, 86, 87], a feature that may alleviate the tension with observational constraints. Importantly, these conclusions do not depend on the detailed shape functions in our examples (i.e., the specific perturbation profiles entering $(\nabla\bar{r})^{2}$); only their relative magnitudes and signs matter [29].
While the qualitative implications are broadly model-independent, a quantitative assessment of both $\Delta\delta$ and $M_{\text{gap}}$ for radiation-fluid collapse demands dedicated semiclassical numerical simulations—a central open challenge for confronting theory with observation.
Acknowledgements.
We thank David Garfinkle, Viqar Husain, Gabor Kunstatter, and Edward Witten for useful discussions. In particular, we thank Roberto Emparan, Carsten Gundlach, and Gustavo J. Turiaci for reading the draft and providing valuable feedback that clarified various subtleties. CHW also thanks Gustavo J. Turiaci for numerous discussions that significantly improved this work. He further acknowledges an audience member at the 39th Pacific Coast Gravity Meeting, whose suggestion inspired this project, and is deeply indebted to Gary Horowitz and Jiuci Xu for their early collaboration and many insightful discussions. CHW is supported by the University of Washington and DOE Award DE-SC0011637. MT is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 852386). MT is also supported by the Emmy Noether Fellowship program at the Perimeter Institute for Theoretical Physics.